you rent a car and are offered 2 payments options. you can pay $25 a day plus 15 c a mile (option a) or you can pay $10 a day plus 40 cents a mile (option b) for what amount of daily miles will option a be the cheaper plan?
Answer by addingup(3677)

25+0.15x = 10+0.40x
15 = 0.25x
x = 15/0.25 = 60
if you drive more than 60 hours plan a is cheaper.
:
Check it out, take an arbitrary number, let's say you drove 80 hours:
a) 25+0.15(80) = $37
b) 10+0.40(80) = $42
you can see that b is more expensive over 60 hours